# Pinwheel Scheduling with Real Periods

### v3

@@ Abstract
For a sequence of tasks, each with a positive integer period, the pinwheel scheduling problem involves finding a valid schedule in the sense that the schedule performs one task per day and each task is performed at least once every consecutive days of its period. It had been conjectured by Chan and Chin in 1993 that there exists a valid schedule for any sequence of tasks with density, the sum of the reciprocals of each period, at most $\frac{5}{6}$. Recently, Kawamura settled this conjecture affirmatively. In this paper we consider an extended version with real periods proposed by Kawamura, in which a valid schedule must perform each task $i$ having a real period $a_{i}$ at least $l$ times in any consecutive $\lceil la_{i}\rceil$ days for all positive integer $l$. We show that any sequence of tasks such that the periods take three distinct real values and the density is at most $\frac{5}{6}$ admits a valid schedule. We hereby conjecture that the conjecture of Chan and Chin is true also for real periods.
00footnotetext: We are grateful to Akitoshi Kawamura for useful discussions. This work was supported by JSPS KAKENHI Grant Numbers 25K14990, 20K11689, and 20K11676.
@@ 1 Introduction
In the pinwheel scheduling problem [HMR+89], we are given a sequence of tasks, each with a positive integer period, and want to find a schedule that performs one task per day such that each task is performed at least once every consecutive days of its period. Let $\mathbb{N}$ and $\mathbb{Z}$ be the set of all positive integers and the set of all integers, respectively, and let $k]=\{i\;|\;i\in\mathbb{N},1\leq i\leq k\}$. Formally, an integer-valued instance of the pinwheel schedule problem is a sequence $A=(a_{i})_{i\in[k]}\in\mathbb{N}^{k}$, where $a_{i}$ is the period of task $i$. The goal is to find a schedule $S:\mathbb{Z}\to[k]$, which indicates that task $S(t)$ is performed on day $t$, such that the following condition is satisfied for each task $i\in[k]$:
@@
Such a schedule is said to be valid for instance $A$. In particular, when a schedule $S$ is periodic, that is, there exists $p\in\mathbb{N}$ such that $S(t)=S(t+p)$ holds for all $t\in\mathbb{Z}$, we represent the schedule by a sequence of $|S(0)S(1)\cdots S(p-1)|$. We say that an instance is schedulable if there exists a valid schedule for the instance. For example, instance $(2,4,4)$ is schedulable, since schedule $|1213|$ is valid for it.
The density of instance $A=(a_{i})_{i\in[k]}$ is defined as $D(A)=\sum_{i\in[k]}\frac{1}{a_{i}}$. Intuitively, the density of an instance means how many tasks a valid schedule for the instance has to perform on average per day. Obviously, $D(A)\leq 1$ is necessary for instance $A$ to admit a valid schedule. (See that the density of instance $(2,4,4)$ above is 1.) However, this is not sufficient. Chan and Chin conjectured as follows:
@@
If an integer-valued instance $A$ satisfies $D(A)\leq\frac{5}{6}$, then $A$ is schedulable.
@@
If an integer-valued instance $A$ satisfies $D(A)\leq\frac{5}{6}$, then $A$ is schedulable.
@@
In other words, this condition is that for each $l\in\mathbb{N}$, task $i$ is performed at least $l$ times in any consecutive $\lceil la_{i}\rceil$ days. Please see that when $a_{i}$ is an integer, the condition is equivalent to that of the original integer-valued problem. For example, a task with period $\frac{7}{2}$ has to be performed at least once every 4 days and at least twice every 7 days. Schedule $|1112112|$ is valid for instance $(2,\frac{7}{2})$.
@@
If a real-valued instance $A$ satisfies $D(A)\leq\frac{5}{6}$, then $A$ is schedulable.
@@
If a real-valued instance $A$ consists of three distinct values and satisfies $D(A)\leq\frac{5}{6}$, then $A$ is schedulable.
@@ 1.1 Previous Results
If an integer-valued instance $A$ consists of three distinct values and satisfies $D(A)\leq\frac{5}{6}$, then $A$ is schedulable.
@@
In the schedule, for each $i=1,2,$ and $3$, identify all the tasks with period $c_{i}$ and rename them as task $i$. If the resulting schedule were valid for instance $(\frac{c_{1}}{q_{1}},\frac{c_{2}}{q_{2}},\frac{c_{3}}{q_{3}})$, this would give a solution to the problem with rational periods. However, this attempt fails. We have a counterexample: instance
@@
Gąsieniec et al. partially solved Conjecture 1 for the case of 12 or fewer tasks [GSW22]. The authors provided a Pareto surface of the whole set of such instances, where a Pareto surface $C$ for a set of pinwheel scheduling instances $I$ is an inclusion minimal set of pairs of an instance with minimum periods and a schedule that is valid for the instance such that, for every instance $A$ in $I$, $C$ includes a valid schedule for $A$. We will also consider this point of view in Sections 2 and 3.
@@
If an integer-valued instance $A=(a_{i})_{i\in[k]}$ satisfies $D(A)\leq 1-\frac{a_{1}-1}{a_{1}(a_{1}+1)}$, then $A$ is schedulable.
@@
If a real-valued instance $A$ contains at most two periods and satisfies $D(A)\leq 1$, then $A$ is schedulable.
@@
If a real-valued instance $A=(a_{i})_{i\in[k]}$ satisfies $D(A)\leq 1-\frac{1+\ln 2}{1+\sqrt{a_{1}}}-\frac{3}{2a_{1}}$, then $A$ is schedulable.
@@ 2 A Warm-Up: Case of Two Distinct Real Values
If a real-valued instance $A$ consists of two distinct values and satisfies $D(A)\leq\frac{5}{6}$, then $A$ is schedulable.
@@
Every real-valued instance $A=(a_{1},\frac{1}{\frac{5}{6}-\frac{1}{a_{1}}})$ with $\frac{6}{5}<a_{1}<\frac{12}{5}$ satisfies at least one of the following conditions and the schedule $S$ defined there is valid for instance $A$:
@@
We denote by $P_{i}(y,l)$ the proposition that for each $m\in\mathbb{Z}$, there exist at least $l$ values of $t\in[m,m+y)\cap\mathbb{Z}$ such that $S(t)=i$, which means that schedule $S$ performs task $i$ at least $l$ times in any consecutive $y$ days. The schedulability of task 1 in instance $(\frac{6}{5},6)$ is represented as $P_{1}(\lceil l\cdot\frac{6}{5}\rceil,l)$ for all $l\in\mathbb{N}$. We are going to prove this.
For each $l=1,2,\ldots,5$, we have $\lceil l\cdot\frac{6}{5}\rceil=l+1$. According to (1), $S(t)=1$ if $t\equiv 0,1,2,3,4$ (mod 6)($\text${mod }6). It is seen that there are $l$ days when task 1 is performed every consecutive $l+1$ days. That is, $P_{1}(\lceil l\cdot\frac{6}{5}\rceil,l)$ is true.
For $l\geq 6$, let $q$ be the quotient and $r$ be the remainder when $l$ is divided by 5. We have $\lceil l\cdot\frac{6}{5}\rceil=\lceil(5q+r)\cdot\frac{6}{5}\rceil=6q+\lceil r\cdot\frac{6}{5}\rceil$. Using what we have already shown, we obtain that: Task 1 is done at least $r$ times in any consecutive $\lceil r\cdot\frac{6}{5}\rceil$ days by $P_{1}(\lceil z\cdot\frac{6}{5}\rceil,z)$ for $z=1,2,3,4$, while it is done at least $5q$ times in any consecutive $6q$ days by $P_{1}(\lceil 5\cdot\frac{6}{5}\rceil,5)$. Therefore, task 1 is done at least $5q+r=l$ times in any consecutive $6q+\lceil r\cdot\frac{6}{5}\rceil$ days, which implies $P_{1}(\lceil l\cdot\frac{6}{5}\rceil,l)$. ∎
@@ 3 Proof of Theorem 2
Every real-valued instance $A=(a_{1},a_{2},\frac{1}{\frac{5}{6}-\frac{1}{a_{1}}-\frac{1}{a_{2}}})$ with $a_{1}<a_{2}$, $\frac{5}{6}<\frac{1}{a_{1}}+\frac{2}{a_{2}}$, and $\frac{1}{a_{1}}+\frac{1}{a_{2}}<\frac{5}{6}$ satisfies at least one of the following conditions and the schedule $S$ defined there is valid for instance $A$:
@@
According to (7), for any $l\in\mathbb{N}$, the maximum time span in which task 1 is performed exactly (l1)(l-1) times is from day $\lceil ma_{1}\rceil$ to day $(\lceil(m+l)a_{1}\rceil-2)$ for some integer $m$, and its length is $(\lceil(m+l)a_{1}\rceil-\lceil ma_{1}\rceil-1)$ days. (i) If $la_{1}$ is an integer, we evaluate $\lceil(m+l)a_{1}\rceil-\lceil ma_{1}\rceil-1=ma_{1}+\lceil la_{1}\rceil-ma_{1}-1=la_{1}-1=\lceil la_{1}\rceil-1$. (ii) Otherwise, we get $\lceil(m+l)a_{1}\rceil-\lceil ma_{1}\rceil-1\leq(\lceil ma_{1}\rceil+\lceil la_{1}\rceil)-\lceil ma_{1}\rceil-1=\lceil la_{1}\rceil-1$. In any case, we observe that task 1 is done at least $l$ times in any consecutive $\lceil la_{1}\rceil$ days.
We finally check the validity of tasks 2 and 3 together. Set $d=0$ for task 2 and $\frac{1}{2}$ for task 3. For any $l\in\mathbb{N}$, the maximum time span in which the task is performed exactly (l1)(l-1) times is from day $(\lfloor(m+d)a_{2}^{\prime}\rfloor+1)$ to day $(\lfloor(m+d+l)a_{2}^{\prime}\rfloor-1)$ for some integer $m$, and its length is $(\lfloor(m+d+l)a_{2}^{\prime}\rfloor-\lfloor(m+d)a_{2}^{\prime}\rfloor-1)$ days. (i) If $la_{2}^{\prime}$ is an integer, we obtain $\lfloor(m+d+l)a_{2}^{\prime}\rfloor-\lfloor(m+d)a_{2}^{\prime}\rfloor-1=\lfloor(m+d)a_{2}^{\prime}\rfloor+la_{2}^{\prime}-\lfloor(m+d)a_{2}^{\prime}\rfloor-1=la_{2}^{\prime}-1=\lceil la_{2}^{\prime}\rceil-1$. (ii) Otherwise, we get $\lfloor(m+d+l)a_{2}^{\prime}\rfloor-\lfloor(m+d)a_{2}^{\prime}\rfloor-1\leq(\lfloor(m+d)a_{2}^{\prime}\rfloor+\lfloor la_{2}^{\prime}\rfloor+1)-\lfloor(m+d)a_{2}^{\prime}\rfloor-1=\lfloor la_{2}^{\prime}\rfloor=\lceil la_{2}^{\prime}\rceil-1$. In any case, we see that the task is done at least $l$ times in any consecutive $\lceil la_{2}^{\prime}\rceil$ days. ∎
@@
Fix $\varepsilon>0$ arbitrarily. As in the proof of Lemma 4, we denote by $P_{i}(y,l)$ the proposition that for each $m\in\mathbb{Z}$, there exist at least $l$ values of $t\in[m,m+y)\cap\mathbb{Z}$ such that $S(t)=i$, which means that schedule $S$ performs task $i$ at least once every consecutive $y$ days. Clearly, for $y\leq y^{\prime}$ and $l\geq l^{\prime}$, $P_{i}(y,l)\Rightarrow P_{i}(y^{\prime},l^{\prime})$ holds true. In addition, if $P_{i}(y,l)$ and $P_{i}(y^{\prime},l^{\prime})$ hold true, then so does $P_{i}(y+y^{\prime},l+l^{\prime})$.
@@
For $l=1$, (8) says that there is at least one day $t$ satisfying $t\equiv 2,5$ (mod 9)($\text${mod }9) in any consecutive $5l+1=6$ days, thus $P_{2}(6,1)\Rightarrow P_{2}(\lceil 1\cdot(5+\varepsilon)\rceil,1)$. For $l=2$, since $\lceil 2\cdot(5+\varepsilon)\rceil\geq 5\cdot 2+1=11>1\cdot 9$, $P_{2}(9,2)\Rightarrow P_{2}(\lceil 2\cdot(5+\varepsilon)\rceil,2)$ holds. For $l=3$, it follows that $\lceil 3\cdot(5+\varepsilon)\rceil\geq 5\cdot 3+1=16=1\cdot 9+7$. Since $P_{2}(6,1)\Rightarrow P_{2}(7,1)$ and $P_{2}(9,2)$, we have $P_{2}(9+7,2+1)\Rightarrow P_{2}(\lceil 3\cdot(5+\varepsilon)\rceil,3)$. For $l=4$, $\lceil 4\cdot(5+\varepsilon)\rceil\geq 5\cdot 4+1=21>2\cdot 9$ holds. Using $P_{2}(9,2)$, we derive $P_{2}(2\cdot 9,2\cdot 2)\Rightarrow P_{2}(\lceil 4\cdot(5+\varepsilon)\rceil,4)$. For $l=5$, we have $\lceil 5\cdot(5+\varepsilon)\rceil\geq 5\cdot 5+1=26=2\cdot 9+8$. From $P_{2}(6,1)\Rightarrow P_{2}(8,1)$ and $P_{2}(9,2)$, we get $P_{2}(2\cdot 9+8,2\cdot 2+1)\Rightarrow P_{2}(\lceil 5\cdot(5+\varepsilon)\rceil,5)$. For $l=6$, $\lceil 6\cdot(5+\varepsilon)\rceil\geq 5\cdot 6+1=31>3\cdot 9$ holds. We have $P_{2}(3\cdot 9,3\cdot 2)\Rightarrow P_{2}(\lceil 6\cdot(5+\varepsilon)\rceil,6)$ by $P_{2}(9,2)$. Finally, for $l\geq 7$, we have $\lceil l\cdot(5+\varepsilon)\rceil\geq 5l+1\geq 9\cdot\lfloor\frac{5l+1}{9}\rfloor$. It also holds true that $2\cdot\lfloor\frac{5l+1}{9}\rfloor\geq l$, since $2\cdot\lfloor\frac{5l+1}{9}\rfloor-l>2\cdot(\frac{5l+1}{9}-1)-l=\frac{l-7}{9}\geq 0$. Applying $P_{2}(9,2)$, we conclude $P_{2}(9\cdot\lfloor\frac{5l+1}{9}\rfloor,2\cdot\lfloor\frac{5l+1}{9}\rfloor)\Rightarrow P_{2}(\lceil l\cdot(5+\varepsilon)\rceil,l)$. ∎

### v4

@@ Abstract
For a sequence of tasks, each with a positive integer period, the pinwheel scheduling problem involves finding a valid schedule in the sense that the schedule performs one task per day and each task is performed at least once every consecutive days of its period. It had been conjectured by Chan and Chin in 1993 that there exists a valid schedule for any sequence of tasks with density, the sum of the reciprocals of each period, at most $\frac{5}{6}$. Recently, Kawamura settled this conjecture affirmatively. In this paper we consider an extended version with real periods proposed by Kawamura, in which a valid schedule must perform each task $i$ having a real period $a_{i}$ at least $l$ times in any $\lceil la_{i}\rceil$ consecutive days for all positive integer $l$. We show that any sequence of tasks such that the periods take three distinct real values and the density is at most $\frac{5}{6}$ admits a valid schedule. We hereby conjecture that the conjecture of Chan and Chin is true also for real periods.
00footnotetext: We are grateful to Akitoshi Kawamura for useful discussions. This work was supported by JSPS KAKENHI Grant Numbers 25K14990, 20K11689, and 20K11676, and also by the Research Institute for Mathematical Sciences (RIMS), an International Joint Usage/Research Center located in Kyoto University.
@@ 1 Introduction
In the pinwheel scheduling problem [HMR+89], we are given a sequence of tasks, each with a positive integer period, and want to find a schedule that performs one task per day such that each task is performed at least once every consecutive days of its period. Let $\mathbb{N}$ and $\mathbb{Z}$ be the set of all positive integers and the set of all integers, respectively, and let $k]=\{i\;|\;i\in\mathbb{N},1\leq i\leq k\}$. Formally, an integer-valued instance of the pinwheel schedule problem is a sequence $A=(a_{i})_{i\in[k]}\in\mathbb{N}^{k}$, where $a_{i}$ is the period of task ii. The goal is to find a schedule $S:\mathbb{Z}\to[k]$, which indicates that task $S(t)$ is performed on day $t$, such that the following condition is satisfied for each task $i\in[k]$:
@@
Such a schedule is said to be valid for instance AA. In particular, when a schedule $S$ is periodic, that is, there exists $p\in\mathbb{N}$ such that $S(t)=S(t+p)$ holds for all $t\in\mathbb{Z}$, we represent the schedule by a sequence of $|S(0)S(1)\cdots S(p-1)|$. We say that an instance is schedulable if there exists a valid schedule for the instance. For example, instance $(2,4,4)$ is schedulable, since schedule $|1213|$ is valid for it.
The density of instance $A=(a_{i})_{i\in[k]}$ is defined as $D(A)=\sum_{i\in[k]}\frac{1}{a_{i}}$. Intuitively, the density of an instance means how many tasks a valid schedule for the instance has to perform on average per day. Obviously, $D(A)\leq 1$ is necessary for instance AA to admit a valid schedule. (See that the density of instance $(2,4,4)$ above is 1.) However, this is not sufficient. Chan and Chin conjectured as follows:
@@
If an integer-valued instance AA satisfies $D(A)\leq\frac{5}{6}$, then AA is schedulable.
@@
If an integer-valued instance AA satisfies $D(A)\leq\frac{5}{6}$, then AA is schedulable.
@@
In other words, this condition is that for each $l\in\mathbb{N}$, task $i$ is performed at least $l$ times in any $\lceil la_{i}\rceil$ consecutive days. Please see that when $a_{i}$ is an integer, the condition is equivalent to that of the original integer-valued problem. For example, a task with period $\frac{7}{2}$ has to be performed at least once every 4 days and at least twice every 7 days. Schedule $|1112112|$ is valid for instance $(2,\frac{7}{2})$.
@@
If a real-valued instance AA satisfies $D(A)\leq\frac{5}{6}$, then AA is schedulable.
@@
If a real-valued instance AA consists of three distinct values and satisfies $D(A)\leq\frac{5}{6}$, then AA is schedulable.
@@ 1.1 Previous Results
If an integer-valued instance AA consists of three distinct values and satisfies $D(A)\leq\frac{5}{6}$, then AA is schedulable.
@@
In the schedule, for each $i=1,2,$ and 33, identify all the tasks with period $c_{i}$ and rename them as task ii. If the resulting schedule were valid for instance $(\frac{c_{1}}{q_{1}},\frac{c_{2}}{q_{2}},\frac{c_{3}}{q_{3}})$, this would give a solution to the problem with rational periods. However, this attempt fails. We have a counterexample: instance
@@
Gąsieniec et al. partially solved Conjecture 1 for the case of 12 or fewer tasks [GSW22]. The authors provided a Pareto surface of the whole set of such instances, where a Pareto surface $C$ for a set of pinwheel scheduling instances II is an inclusion minimal set of pairs of an instance with minimum periods and a schedule that is valid for the instance such that, for every instance AA in II, $C$ includes a valid schedule for AA. We will also consider this point of view in Sections 2 and 3.
@@
If an integer-valued instance $A=(a_{i})_{i\in[k]}$ satisfies $D(A)\leq 1-\frac{a_{1}-1}{a_{1}(a_{1}+1)}$, then AA is schedulable.
@@
If a real-valued instance AA contains at most two periods and satisfies $D(A)\leq 1$, then AA is schedulable.
@@
If a real-valued instance $A=(a_{i})_{i\in[k]}$ satisfies $D(A)\leq 1-\frac{1+\ln 2}{1+\sqrt{a_{1}}}-\frac{3}{2a_{1}}$, then AA is schedulable.
@@ 2 A Warm-Up: Case of Two Distinct Real Values
If a real-valued instance AA consists of two distinct values and satisfies $D(A)\leq\frac{5}{6}$, then AA is schedulable.
@@
Every real-valued instance $A=(a_{1},\frac{1}{\frac{5}{6}-\frac{1}{a_{1}}})$ with $\frac{6}{5}<a_{1}<\frac{12}{5}$ satisfies at least one of the following conditions and the schedule $S$ defined there is valid for instance AA:
@@
We denote by $P_{i}(y,l)$ the proposition that for each $m\in\mathbb{Z}$, there exist at least $l$ values of $t\in[m,m+y)\cap\mathbb{Z}$ such that $S(t)=i$, which means that schedule $S$ performs task $i$ at least $l$ times in any $y$ consecutive days. The schedulability of task 1 in instance $(\frac{6}{5},6)$ is represented as $P_{1}(\lceil l\cdot\frac{6}{5}\rceil,l)$ for all $l\in\mathbb{N}$. We are going to prove this.
For each $l=1,2,\ldots,5$, we have $\lceil l\cdot\frac{6}{5}\rceil=l+1$. According to (1), $S(t)=1$ if $t\equiv 0,1,2,3,4$ (mod 6)($\text${mod }6). It is seen that there are $l$ days when task 1 is performed every $l+1$ consecutive days. That is, $P_{1}(\lceil l\cdot\frac{6}{5}\rceil,l)$ is true.
For $l\geq 6$, let $q$ be the quotient and $r$ be the remainder when $l$ is divided by 5. We have $\lceil l\cdot\frac{6}{5}\rceil=\lceil(5q+r)\cdot\frac{6}{5}\rceil=6q+\lceil r\cdot\frac{6}{5}\rceil$. Using what we have already shown, we obtain that: Task 1 is done at least $r$ times in any $\lceil r\cdot\frac{6}{5}\rceil$ consecutive days by $P_{1}(\lceil z\cdot\frac{6}{5}\rceil,z)$ for $z=1,2,3,4$, while it is done at least $5q$ times in any $6q$ consecutive days by $P_{1}(\lceil 5\cdot\frac{6}{5}\rceil,5)$. Therefore, task 1 is done at least $5q+r=l$ times in any $6q+\lceil r\cdot\frac{6}{5}\rceil$ consecutive days, which implies $P_{1}(\lceil l\cdot\frac{6}{5}\rceil,l)$. ∎
@@ 3 Proof of Theorem 2
Every real-valued instance $A=(a_{1},a_{2},\frac{1}{\frac{5}{6}-\frac{1}{a_{1}}-\frac{1}{a_{2}}})$ with $a_{1}<a_{2}$, $\frac{5}{6}<\frac{1}{a_{1}}+\frac{2}{a_{2}}$, and $\frac{1}{a_{1}}+\frac{1}{a_{2}}<\frac{5}{6}$ satisfies at least one of the following conditions and the schedule $S$ defined there is valid for instance AA:
@@
According to (7), for any $l\in\mathbb{N}$, the maximum time span in which task 1 is performed exactly (l1)(l-1) times is from day $\lceil ma_{1}\rceil$ to day $(\lceil(m+l)a_{1}\rceil-2)$ for some integer $m$, and its length is $(\lceil(m+l)a_{1}\rceil-\lceil ma_{1}\rceil-1)$ days. (i) If $la_{1}$ is an integer, we evaluate $\lceil(m+l)a_{1}\rceil-\lceil ma_{1}\rceil-1=ma_{1}+\lceil la_{1}\rceil-ma_{1}-1=la_{1}-1=\lceil la_{1}\rceil-1$. (ii) Otherwise, we get $\lceil(m+l)a_{1}\rceil-\lceil ma_{1}\rceil-1\leq(\lceil ma_{1}\rceil+\lceil la_{1}\rceil)-\lceil ma_{1}\rceil-1=\lceil la_{1}\rceil-1$. In any case, we observe that task 1 is done at least $l$ times in any $\lceil la_{1}\rceil$ consecutive days.
We finally check the validity of tasks 2 and 3 together. Set $d=0$ for task 2 and $\frac{1}{2}$ for task 3. For any $l\in\mathbb{N}$, the maximum time span in which the task is performed exactly (l1)(l-1) times is from day $(\lfloor(m+d)a_{2}^{\prime}\rfloor+1)$ to day $(\lfloor(m+d+l)a_{2}^{\prime}\rfloor-1)$ for some integer $m$, and its length is $(\lfloor(m+d+l)a_{2}^{\prime}\rfloor-\lfloor(m+d)a_{2}^{\prime}\rfloor-1)$ days. (i) If $la_{2}^{\prime}$ is an integer, we obtain $\lfloor(m+d+l)a_{2}^{\prime}\rfloor-\lfloor(m+d)a_{2}^{\prime}\rfloor-1=\lfloor(m+d)a_{2}^{\prime}\rfloor+la_{2}^{\prime}-\lfloor(m+d)a_{2}^{\prime}\rfloor-1=la_{2}^{\prime}-1=\lceil la_{2}^{\prime}\rceil-1$. (ii) Otherwise, we get $\lfloor(m+d+l)a_{2}^{\prime}\rfloor-\lfloor(m+d)a_{2}^{\prime}\rfloor-1\leq(\lfloor(m+d)a_{2}^{\prime}\rfloor+\lfloor la_{2}^{\prime}\rfloor+1)-\lfloor(m+d)a_{2}^{\prime}\rfloor-1=\lfloor la_{2}^{\prime}\rfloor=\lceil la_{2}^{\prime}\rceil-1$. In any case, we see that the task is done at least $l$ times in any $\lceil la_{2}^{\prime}\rceil$ consecutive days. ∎
@@
Fix $\varepsilon>0$ arbitrarily. As in the proof of Lemma 4, we denote by $P_{i}(y,l)$ the proposition that for each $m\in\mathbb{Z}$, there exist at least $l$ values of $t\in[m,m+y)\cap\mathbb{Z}$ such that $S(t)=i$, which means that schedule $S$ performs task $i$ at least once every $y$ consecutive days. Clearly, for $y\leq y^{\prime}$ and $l\geq l^{\prime}$, $P_{i}(y,l)\Rightarrow P_{i}(y^{\prime},l^{\prime})$ holds true. In addition, if $P_{i}(y,l)$ and $P_{i}(y^{\prime},l^{\prime})$ hold true, then so does $P_{i}(y+y^{\prime},l+l^{\prime})$.
@@
For $l=1$, (8) says that there is at least one day $t$ satisfying $t\equiv 2,5$ (mod 9)($\text${mod }9) in any $5l+1=6$ consecutive days, thus $P_{2}(6,1)\Rightarrow P_{2}(\lceil 1\cdot(5+\varepsilon)\rceil,1)$. For $l=2$, since $\lceil 2\cdot(5+\varepsilon)\rceil\geq 5\cdot 2+1=11>1\cdot 9$, $P_{2}(9,2)\Rightarrow P_{2}(\lceil 2\cdot(5+\varepsilon)\rceil,2)$ holds. For $l=3$, it follows that $\lceil 3\cdot(5+\varepsilon)\rceil\geq 5\cdot 3+1=16=1\cdot 9+7$. Since $P_{2}(6,1)\Rightarrow P_{2}(7,1)$ and $P_{2}(9,2)$, we have $P_{2}(9+7,2+1)\Rightarrow P_{2}(\lceil 3\cdot(5+\varepsilon)\rceil,3)$. For $l=4$, $\lceil 4\cdot(5+\varepsilon)\rceil\geq 5\cdot 4+1=21>2\cdot 9$ holds. Using $P_{2}(9,2)$, we derive $P_{2}(2\cdot 9,2\cdot 2)\Rightarrow P_{2}(\lceil 4\cdot(5+\varepsilon)\rceil,4)$. For $l=5$, we have $\lceil 5\cdot(5+\varepsilon)\rceil\geq 5\cdot 5+1=26=2\cdot 9+8$. From $P_{2}(6,1)\Rightarrow P_{2}(8,1)$ and $P_{2}(9,2)$, we get $P_{2}(2\cdot 9+8,2\cdot 2+1)\Rightarrow P_{2}(\lceil 5\cdot(5+\varepsilon)\rceil,5)$. For $l=6$, $\lceil 6\cdot(5+\varepsilon)\rceil\geq 5\cdot 6+1=31>3\cdot 9$ holds. We have $P_{2}(3\cdot 9,3\cdot 2)\Rightarrow P_{2}(\lceil 6\cdot(5+\varepsilon)\rceil,6)$ by $P_{2}(9,2)$. Finally, for $l\geq 7$, we have $\lceil l\cdot(5+\varepsilon)\rceil\geq 5l+1\geq 9\cdot\lfloor\frac{5l+1}{9}\rfloor$. It also holds true that $2\cdot\lfloor\frac{5l+1}{9}\rfloor\geq l$, since $2\cdot\lfloor\frac{5l+1}{9}\rfloor-l>2\cdot(\frac{5l+1}{9}-1)-l=\frac{l-7}{9}\geq 0$. Applying $P_{2}(9,2)$, we conclude $P_{2}(9\cdot\lfloor\frac{5l+1}{9}\rfloor,2\cdot\lfloor\frac{5l+1}{9}\rfloor)\Rightarrow P_{2}(\lceil l\cdot(5+\varepsilon)\rceil,l)$. ∎